OpenStudy (anonymous):

an unconditional acceptance into a graduate program at a university will be given to students whose GMAT score plus 100 times the undergraduate grade point average is at least 1000. Robbin's GMAT score was 750. What must her grade point average be in order to be unconditionally accepted into the program?

OpenStudy (anonymous):

6-5X>3-4X

OpenStudy (anonymous):

(GMAT score + 400) X (Undergrad Grade Point Average) >= 1800 1200×UndergradGradePointAverage≥1800 Average≥1800/1200=3/2=1.5 Bigger or equal to 1.5